OpenStudy (anonymous):

Problem: A curve C is the locus of points in the xy-plane which are equidistant from the point (0,7.25 ) and the line (y=-7.25). Identify the curve?

OpenStudy (tkhunny):

This should scream PARABOLA!!!! It is the fundamental definition. (0,7.25) is the Focus y = - 7.25 is the Directrix The vertex just might be at the Origin!

OpenStudy (anonymous):

may be it has some sense but i cant understand why.

OpenStudy (tkhunny):

So, it didn't scream? Oh, well. Maybe we can just do the algebra. equidistant from the point (0,7.25 ) the line (y=-7.25) Any old point might be described as (a,b). How far is (a,b) from y = -7.25 = -29/4?

OpenStudy (anonymous):

well, since point have coordinate (0, 7.25) then distance is \[D=\sqrt{a^2+(b-7.25)^2}\]

OpenStudy (tkhunny):

Good, you answered my next question. That is the distance from the point. We certainly need that. How about the other one. How far is (a,b) from y = -29/4?

OpenStudy (anonymous):

well no clue how to find distance from (a,b) to line, since it can have different values.

OpenStudy (tkhunny):

b - (-29/4) -- On need consider only the y-value

OpenStudy (anonymous):

how it give us a parabola then?

OpenStudy (tkhunny):

If you believe that last piece, we are set. EQUIDISTANT \(b + 7.25 = \sqrt{a^{2} + (b-7.25)^{2}}\) Now, there's a little algebra to go. First, and foremost, we're talking about distances, so everything is positive and there is no harm in squaring things. \((b + 7.25)^{2} = a^{2} + (b-7.25)^{2}\) Is it looking better, yet?

OpenStudy (anonymous):

now l start to geting something about it.

OpenStudy (tkhunny):

Okay, expand and simplify and you should see it!

OpenStudy (anonymous):

thx you.

OpenStudy (tkhunny):

BTW - Excellent work coming up with that distance formula. This suggests you are paying attention. Medal for you!